# Multi-Step Knowledge Interaction Analysis via Rank-2 Subspace Disentanglement

### v1

@@ Abstract
Natural Language Explanations (NLEs) describe how Large Language Models (LLMs) make decisions, drawing on both external Context Knowledge (CK) and Parametric Knowledge (PK) stored in model weights. Understanding their interaction is key to assessing the grounding of NLEs, yet it remains underexplored. Prior work has largely examined only single-step generation – typically the final answer, and has modelled PK–CK interaction only as a binary choice in a rank-1 subspace. This overlooks richer forms of interaction, such as complementary or supportive knowledge. We propose a novel rank-2 projection subspace that disentangles PK and CK contributions more accurately and use it for the first multi-step analysis of knowledge interactions across longer NLE sequences. Experiments on four QA datasets and three open-weight instruction-tuned LLMs show that diverse knowledge interactions are poorly represented in a rank-1 subspace but are effectively captured in our rank-2 formulation. Our multi-step analysis reveals that hallucinated NLEs align strongly with the PK direction, context-faithful ones balance PK and CK, and Chain-of-Thought prompting for NLEs shifts generated NLEs toward CK by reducing PK reliance. This work provides the first framework for systematic studies of multi-step knowledge interactions in LLMs through a richer rank-2 subspace disentanglement. [^1] Code and data: https://github.com/copenlu/pk-ck-knowledge-disentanglement
Figure 1: Llama-3.1-8B-Instruct model combines parametric (green) and contextual (red) knowledge to generate NLEs. Projection onto a learned low-rank subspace $\mathbf{P}$ disentangles their contributions – rank-1 discards richer interactions, while rank-2 separates complementary and conflicting components.
@@ 1 Introduction
Large Language Models (LLMs) are employed to generate Natural Language Explanations (NLEs) in a human-readable format, illustrating the underlying decision-making process for predictions in complex reasoning tasks such as Claim Verification (CV) and Question Answering (QA). These NLEs are valuable because they can reveal the utilization of external context and the knowledge stored in model parameters. Consider the Fig. 1 example from a QA task (Cheng et al., 2024), where Llama-3.1-8B-Instruct Meta-Team (2024) generates the NLE by utilizing both Context Knowledge (CK) and Parametric Knowledge (PK), to explain the underlying decision-making process for the final answer prediction. For some tasks, such as CV, we assume the decision-making process and, in turn, the NLE, will rely more on evidence (CK) Wang and Shu (2023); Tan et al. (2025), and in QA tasks with misleading external context, we assume they will rely on the PK. On the other hand, Chain-of-Thought (CoT) prompting (Wei et al., 2022), a widely adopted reasoning methodology, explicitly elicits NLEs of the intermediate reasoning steps and is hypothesized to influence how LLMs integrate PK and CK (Cheng et al., 2024; Su et al., 2024; Tao et al., 2025), potentially improving contextual grounding. However, it remains unclear what the learned PK-CK interaction dynamics in generating the NLEs are, necessitating a multi-step analysis of PK-CK interaction across longer NLE sequences.
@@ 2 Related Work
Parametric vs. contextual knowledge in LMs. Prior work has extensively examined what LMs store as parametric knowledge (PK) and how they integrate context knowledge (CK) during inference (Petroni et al., 2019; Jiang et al., 2020; Roberts et al., 2020; Hagström et al., 2025; Marjanovic et al., 2024; Yu et al., 2024). Early studies showed that LMs act as factual knowledge bases, with knowledge retrievable via prompts (Petroni et al., 2019), while probing analyses disentangled what is encoded in the parameters versus what is acquired from context (Brown et al., 2020; Tenney et al., 2019; Bi et al., 2025).
@@
Recent research has explored the interplay and conflict between PK and CK. Cheng et al. (2024) revealed PK suppression under various knowledge interactions, and Xu et al. (2024) categorized conflict types (context–memory, inter-context, intra-memory) to study resulting behavioral shifts. Enhancing PK–CK balance has been pursued through improved pretraining or fine-tuning (Zhang et al., 2024), context-aware representation interventions (Yuan et al., 2025b), contrastive decoding (Zhao et al., 2024), and lightweight steering methods that increase context sensitivity without weight updates (Wang et al., 2025). Our work extends this literature by providing a geometric framework that accurately models PK–CK interactions in a rank-2 subspace and enables multi-step analysis of how both knowledge sources jointly shape Natural Language Explanations (NLEs).
NLEs and explanation-guided reasoning. Natural Language Explanations (NLEs) Camburu et al. (2018); Atanasova et al. (2020) have been used to expose, supervise, and steer model reasoning in complex tasks. Human-written NLEs improve commonsense QA via explanation-based fine-tuning (Rajani et al., 2019), while Chain-of-Thought (CoT) prompting elicits stepwise reasoning that boosts performance on multi-hop and arithmetic tasks (Wei et al., 2022). Explanations have also been used as few-shot exemplars to enhance in-context learning (Lampinen et al., 2022).
Despite their utility, studies reveal that NLEs often fail to reflect true model reasoning: counterfactual and reconstruction tests expose limited faithfulness (Atanasova et al., 2023; Siegel et al., 2024; Yuan et al., 2025a; Wang and Atanasova, 2025), and CoT explanations can rationalize biased outputs rather than genuine reasoning (Turpin et al., 2023). Our work complements these findings by providing a geometric, token-level analysis of NLE generation, quantifying how parametric and contextual knowledge jointly shape explanation faithfulness and hallucination.
Probing, subspaces, and identifiability. Probing methods uncover low-dimensional structures in language model representations that encode linguistic or factual properties. Hewitt and Manning (2019) showed that syntactic relations are linearly recoverable from word representations, with parse depth and dependency distance captured via geometric transformations. Clark et al. (2019) identified specific BERT attention heads that encode syntactic and semantic cues through linear probes. Elhage et al. (2022) analyzed neuron–feature mappings and proposed that LMs compress information via low-dimensional orthogonal projections, representing dense and sparse features through feature superposition.
Building on this view, Minder et al. (2025) modeled PK–CK conflict as a single controllable direction – a rank-1 projection subspace. However, if PK and CK are superposed within the representation space, a single direction cannot uniquely disentangle their contributions. Motivated by this, we generalize to a rank-2 projection subspace that establishes identifiability of token-level PK and CK components. We futher use this more accurate representation of CK-PK interactions to provide the first multi-step analysis of their interaction dynamics during NLE generation.
@@ 3.1 Task Formulation
Let $\mathcal{Q}$ and $\mathcal{C}$ denote two disjoint sets of queries and contexts, respectively, and consider a QA dataset $\mathcal{D}\subseteq\mathcal{Q}\times\mathcal{C}$. For a question–context pair $(q,c)\in\mathcal{D}$, with $q\in\mathcal{Q}$ and $c\in\mathcal{C}$, a language model $p$ generates both an answer $a$ and an NLE with $n$ tokens as $\mathcal{E}=\{e_{i}\}_{i=1}^{n}$. We distinguish between three forms of answer generation: (i) the parametric answer $a(q,\varepsilon)$, produced by recalling PK (independent of $c$, $\varepsilon$ denotes no context), (ii) the contextual answer $a(q,c)$, obtained by leveraging the provided context while disregarding parametric recall, and (iii) the final predicted answer $a$ combining both the provided context and the PK recall guided by the PK-CK knowledge interaction.
During the sequential generation of $\mathcal{E}$, each token $e_{i}$ is influenced by the interaction between $a(q,\varepsilon)$ and $a(q,c)$. This evolving interaction guides the generation of NLE $\mathcal{E}$, illustrating the decision-making process of the final predicted answer $a$. To analyze this process, we quantify the contribution of parametric knowledge ($\alpha_{i}^{p}$) and contextual knowledge ($\alpha_{i}^{c}$) at each generation step $i$, and we define their difference as
@@ 3.2 Identify Different PK-CK Interactions
To characterize different types of PK–CK interactions guided by individual knowledge contribution, we draw from Minder et al. (2025), who analyze intent-driven answer control, i.e., controlling the model towards a specific answer generation aligned with the intent of following either PK or CK. Let $w_{c}$ (instruction to follow the CK only) and $w_{p}$ (instruction to follow the PK only) denote intents toward predicting $a(q,c)$ and $a(q,\varepsilon)$, respectively, for a given $(q,c)$. The joint intent $w_{b}=\{w_{c},w_{p}\}$ shapes how the model balances contextual and parametric influences. Formally, this intent is encoded within the parameters of $p$, creating a causal relationship between $w$ and the observed $\Delta_{i}$ dynamics. The prompt template for the three intent-driven answer control is described in Tab. 1 in §A.1.2.
Minder et al. (2025) restrict intent to conflicting cases, where $a(q,c)\neq a(q,\varepsilon)$ and the final answer $a$ is determined by either $w_{c}$ or $w_{p}$. We generalize this, following Cheng et al. (2024), to encompass a broader set of interactions:
  • Supportive: $a(q,c)=a(q,\varepsilon)$; PK and CK reinforce the same outcome.
  • Complementary: $a(q,c)\neq a(q,\varepsilon)$ but $a=a(q,c)\land a(q,\varepsilon)$; PK and CK provide non-overlapping but mutually useful contributions.
  • Conflicting: $a(q,c)\neq a(q,\varepsilon)$ and $a$ reflects either $w_{c}$ or $w_{p}$ exclusively.
  • Irrelevant: $a=a(q,\varepsilon)$ while $a(q,\varepsilon)\perp c$; PK dominates, ignoring contextual input.
This taxonomy enables a systematic investigation of how PK and CK interact throughout NLE generation, moving beyond a binary conflict to capture a spectrum of dynamics. We reformulate this intent-driven answer control and $w_{b}$ denote the joint intent of $w_{c}$ and $w_{p}$ that illustrates the intrinsic model behaviour in generating the final answer $a$, considering both the PK and CK governed by the underlying interaction scenario. Unlike in Minder et al. (2025), where $w_{c}$ and $w_{p}$ are represented using one orthonormal direction in a learned rank-1 projection subspace, we represent $w_{c}$ and $w_{p}$ using two orthonormal directions, and $w_{b}\rightarrow w_{c}$ and $w_{b}\rightarrow w_{p}$ indicate the individual contributions of CK and PK, respectively, in generating $a$. Then the joint intent $w_{b}=f(w_{c},w_{p})$ indicates different knowledge interaction scenarios decided by individual contributions from $w_{c}$ and $w_{p}$. We learn this knowledge interaction function using a rank-2 projection subspace.
@@ 3.3 Localize Intent-Guided PK-CK Interactions
Identifying important layers for rank-1 projection subspace. We follow the activation patching-based mechanistic interpretability approach, Patchscope Minder et al. (2025). They construct two minimally different prompts $s$ and $t$ as ‘source’ and ‘target’, with the same $q$ and $c$, only differing by the intent, resulting in two different intended answers. To identify the layers capturing the intent present in $s$, during the forward pass of the target $p(.|t)$, the activation from the hidden state of a particular layer is replaced by the activations from the same layer during the forward pass with the source $p(.|s)$, resulting in $\tilde{p}(.|t)\approx p(.|s)$. Since, they only consider the ‘conflicting’ behaviours in understanding the knowledge intearction in LMs, they construct two patching datasets capturing opposite directions: $\mathcal{D}_{w}^{(c\rightarrow p)}=\{((q,c,w_{c}),a(q,c)),\;((q,c,w_{p}),a(q,\varepsilon))\}$ for encoding the CK direction, and $\mathcal{D}_{w}^{(p\rightarrow c)}=\{((q,c,w_{p}),a(q,\varepsilon)),\;((q,c,w_{c}),a(q,c))\}$ for encoding the PK direction. Layers capturing CK and PK directions, respectively, are selected as:
@@
$\mathbb{L}_{p\rightarrow c}$ $=\{\,l\in[1,L]\mid\tilde{p}(a(q,c)\mid q,c,w_{p})\geq\tau_{p}\,\},$
where $\tau_{c}$, $\tau_{p}$ are hyperparameters (Tab 12 in §A.1.3). For more details, please refer to Minder et al. (2025).
Identifying important layers for rank-2 projection subspace (showing the individual knowledge contribution towards generating $a$, i.e, $w_{b}\rightarrow w_{c}$ and $w_{b}\rightarrow w_{p}$). We again follow Patchscope and prepare two patching datasets $\mathcal{D}_{w}^{(b\rightarrow p)}=\{((q,c,w_{b}),a),\;((q,c,w_{p}),a(q,\varepsilon))\}$ for encoding the contribution of PK interacting with CK in generating $a$, and $\mathcal{D}_{w}^{(b\rightarrow c)}=\{((q,c,w_{b}),a),\;((q,c,w_{c}),a(q,c))\}$ for encoding the contribution of CK interacting with PK in generating $a$. Important layers $\mathbb{L}_{b\rightarrow c}$ and $\mathbb{L}_{b\rightarrow p}$ are selected using Eq. 2.
@@
Theorem 1 (Non-identifiability under rank-1).
@@
Let $a=\langle\vec{\mathbf{v}},\vec{\mathbf{u}}_{CK}\rangle$ and $b=\langle\vec{\mathbf{v}},\vec{\mathbf{u}}_{PK}\rangle$. For any $(c,p)$, choose $(c^{\prime},p^{\prime})=(c+\delta,\,p+\tfrac{a}{b}\delta)$ with $\delta\neq 0$. Then $ac+bp=ac^{\prime}+bp^{\prime}$, so infinitely many $(c^{\prime},p^{\prime})$ yield the same observation. Thus the mapping $(c,p)\mapsto\alpha$ is non-injective, and the individual contributions cannot be disentangled. ∎
Encoding the intent via rank-2 subspace projection. Once we identify important layers for $w_{b}\rightarrow w_{c}$ and $w_{b}\rightarrow w_{p}$, we learn the joint intent function $f$ encoding how individual knowledge contributions $w_{b}\rightarrow w_{c}$ and $w_{b}\rightarrow w_{p}$ are mixed to generate the final answer $a$ using the common layers from $\mathbb{L}_{b\rightarrow c}$ and $\mathbb{L}_{b\rightarrow p}$ and a rank-2 projection subspace $P$ spanned by two orthogonal directions $\vec{\mathbf{u}}\in\mathbb{R}^{d*2}$ as $\mathbf{P}=\vec{\mathbf{u}}(\vec{\mathbf{u}}\vec{\mathbf{u}}^{T})^{-1}\vec{\mathbf{u}}^{T}$. Since we aim to identify individual contributions from the individual directions, following (Minder et al., 2025), we consider those two basis vectors as orthonormal and hence the rank-2 projection subspace $\mathbf{P}$ is reduced to $\mathbf{P}=\vec{\mathbf{u}}\vec{\mathbf{u}}^{T}$, since $\vec{\mathbf{u}}\vec{\mathbf{u}}^{T}=\mathbf{I}$. Considering $\vec{\mathbf{u}}=[\vec{\mathbf{u_{c}}};\vec{\mathbf{u_{p}}}]$, at any sequence step $i$, the hidden representation $\vec{\mathbf{h_{i}}}\in\mathbb{R}^{d}$ for the token $e_{i}$ in the NLE $\mathcal{E}$ can be linearly decomposed as:
@@
Let $c_{i}=\langle\vec{\mathbf{u_{c}}}^{T},\vec{\mathbf{h_{i}}}\rangle$ and $p_{i}=\langle\vec{\mathbf{u_{p}}}^{T},\vec{\mathbf{h_{i}}}\rangle$, then the normalized contribution from CK and PK can be computed as ${\alpha_{i}}^{c}=c_{i}/(c_{i}+p_{i})$ and ${\alpha_{i}}^{p}=p_{i}/(c_{i}+p_{i})$, and ${\alpha_{i}}^{c},{\alpha_{i}}^{p}\in[0,1]$ satisfying the identifiability of individual knowledge contribution (Theorem §1) under rank-2 subspace. The methodology of assigning PK-CK direction to the basis vector $\vec{\mathbf{u}}$ is described in §A.1.1.

### v2

@@ Abstract
Natural Language Explanations (NLEs) describe how Large Language Models (LLMs) make decisions by drawing on external Context Knowledge (CK) and Parametric Knowledge (PK). Understanding the interaction between these sources is key to assessing NLE grounding, yet these dynamics remain underexplored. Prior work has largely focused on i) single-step generation and ii) modelled PK–CK interaction as a binary choice within a rank-1 subspace. This approach overlooks richer interactions and how they unfold over longer generations, such as complementary or supportive knowledge. We propose a novel rank-2 projection subspace that disentangles PK and CK contributions more accurately and use it for the first multi-step analysis of knowledge interactions across longer NLE sequences. Experiments across four QA datasets and three open-weight LLMs demonstrate that while rank-1 subspaces struggle to represent diverse interactions, our rank-2 formulation captures them effectively, highlighting PK alignment for supportive interactions and CK alignment for conflicting ones. Our multi-step analysis reveals among others that hallucinated generations exhibit strong alignment with the PK direction, whereas context-faithful generations maintain a more balanced alignment between PK and CK [^1] Code and data: https://github.com/copenlu/pk-ck-knowledge-disentanglement.
Machine Learning, ICML
Figure 1: (Left) Llama-3.1-8B-Instruct model combines parametric (green) and contextual (orange) knowledge to generate NLEs. (Right) Projection of the answer token to a learned low-rank subspace $\mathbf{P}$ disentangles their contributions – (a) rank-1 fails to disentangle individual knowledge contributions and its orthogonal subspace encodes the correct knowledge, whereas (b) rank-2 successfully encodes the correct knowledge and disentangles individual knowledge contributions.
@@ 1 Introduction
Large Language Models (LLMs) are employed to generate Natural Language Explanations (NLEs) in a human-readable format, illustrating the underlying decision-making process for predictions in complex reasoning tasks such as Claim Verification (CV) and Question Answering (QA). These NLEs are valuable because they can reveal the utilization of external context and the knowledge stored in model parameters. Consider the Figure 1 example from a QA task (Cheng et al., 2024), where Llama-3.1-8B-Instruct (Meta-Team, 2024) generates the NLE by utilizing both Context Knowledge (CK) and Parametric Knowledge (PK), to explain the underlying decision-making process for the final answer prediction. For some tasks, such as CV, we assume the decision-making process and, in turn, the NLE, will rely more on evidence (CK) (Wang and Shu, 2023; Tan et al., 2025), and in QA tasks with misleading external context, we assume they will rely on the PK. On the other hand, Chain-of-Thought (CoT) prompting (Wei et al., 2022), a widely adopted reasoning methodology, explicitly elicits NLEs of the intermediate reasoning steps and is hypothesized to influence how LLMs integrate PK and CK (Cheng et al., 2024; Su et al., 2024; Tao et al., 2025), potentially improving contextual grounding. However, it remains unclear what the learned PK-CK interaction dynamics in generating the NLEs are, necessitating a multi-step analysis of PK-CK interaction across longer NLE sequences.
@@ 2 Related Work
Parametric vs. contextual knowledge in LMs. A large body of work studies what language models encode as parametric knowledge (PK) and how they integrate contextual knowledge (CK) at inference time (Petroni et al., 2019; Jiang et al., 2020; Roberts et al., 2020; Hagström et al., 2025; Marjanovic et al., 2024; Yu et al., 2024). Early work showed that factual knowledge can be elicited directly from model parameters (Petroni et al., 2019), while probing and representation analyses examined how context and memory are encoded and combined across layers (Brown et al., 2020; Tenney et al., 2019; Bi et al., 2025).
@@
More recent studies focus on PK–CK interaction and conflict. Cheng et al. (2024) analyzed suppression effects between PK and CK under different interaction regimes, while Xu et al. (2024) categorized conflict types (context–memory, inter-context, intra-memory) and their behavioral consequences. Several approaches aim to improve PK–CK balance through training or inference-time interventions, including enhanced pretraining or fine-tuning (Zhang et al., 2024), context-aware representation manipulation (Yuan et al., 2025b), contrastive decoding (Zhao et al., 2024), and lightweight steering methods that increase context sensitivity without weight updates (Wang et al., 2025). Our work complements this line by introducing a geometric framework that disentangles PK and CK in a rank-2 subspace and enables multi-step analysis during generation.
Natural language explanations and reasoning. Natural Language Explanations (NLEs) are widely used to expose or guide model reasoning in complex tasks (Camburu et al., 2018; Atanasova et al., 2020). Explanation-based supervision improves QA performance (Rajani et al., 2019), and Chain-of-Thought (CoT) prompting elicits stepwise reasoning that enhances multi-hop and arithmetic reasoning (Wei et al., 2022; Lampinen et al., 2022). However, multiple studies show that NLEs and CoT are often unfaithful to the model’s true decision process, functioning as post-hoc rationalizations rather than causal explanations (Atanasova et al., 2023; Turpin et al., 2023; Siegel et al., 2024; Yuan et al., 2025a; Wang and Atanasova, 2025). Our work addresses this gap by providing a token-level, representation-based analysis of how PK and CK jointly shape NLE generation, linking explanation faithfulness and hallucination to internal knowledge alignment.
Probing, subspaces, and identifiability. Probing methods reveal that linguistic and factual properties are encoded in low-dimensional subspaces of LM representations (Hewitt and Manning, 2019; Clark et al., 2019). The superposition hypothesis further suggests that multiple features are compressed into overlapping linear subspaces (Elhage et al., 2022). Building on this view, Minder et al. (2025) modeled PK–CK conflict using a single controllable direction (rank-1 subspace). However, when PK and CK are jointly encoded, such a representation is non-identifiable and cannot disentangle individual contributions. We generalize this approach to a rank-2 projection subspace, establishing identifiability and enabling fine-grained, multi-step analysis of PK–CK interaction dynamics during generation.
@@ 3.1 Task Formulation
Let $\mathcal{Q}$ and $\mathcal{C}$ denote two disjoint sets of queries and contexts, respectively, and consider a QA dataset $\mathcal{D}\subseteq\mathcal{Q}\times\mathcal{C}$. For a question–context pair $(q,c)\in\mathcal{D}$, with $q\in\mathcal{Q}$ and $c\in\mathcal{C}$, a language model $p$ generates both an answer $a$ and an NLE with $n$ tokens as $\mathcal{E}=\{e_{i}\}_{i=1}^{n}$. We distinguish between three forms of answer generation: (i) the parametric answer $a(q,\phi)$, produced by recalling PK (independent of $c$, $\phi$ denotes no context), (ii) the contextual answer $a(q,c)$, obtained by leveraging the provided context while disregarding parametric recall, and (iii) the final predicted answer $a$ combining both the provided context and the PK recall guided by the PK-CK knowledge interaction.
During the sequential generation of $\mathcal{E}$, each token $e_{i}$ is influenced by the interaction between $a(q,\phi)$ and $a(q,c)$. This evolving interaction guides the generation of NLE $\mathcal{E}$, illustrating the decision-making process of the final predicted answer $a$. To analyze this process, we quantify the contribution of parametric knowledge ($\alpha_{i}^{p}$) and contextual knowledge ($\alpha_{i}^{c}$) at each generation step $i$, and we define their difference as
@@ 3.2 Identify Different PK-CK Interactions
To characterize different types of PK–CK interactions guided by individual knowledge contribution, we draw from Minder et al. (2025), who analyze intent-driven answer control, i.e., controlling the model towards a specific answer generation aligned with the intent of following either PK or CK. Let $w_{c}$ (instruction to follow the CK only) and $w_{p}$ (instruction to follow the PK only) denote intents toward predicting $a(q,c)$ and $a(q,\phi)$, respectively, for a given $(q,c)$.
Minder et al. (2025) restrict intent to conflicting cases, where $a(q,c)\neq a(q,\phi)$ and the final answer $a$ is determined by either $w_{c}$ or $w_{p}$. We generalize this, following Cheng et al. (2024), to encompass a broader set of interactions:
  • Supportive: $a(q,c)=a(q,\phi)=a$; PK and CK reinforce the same outcome, leading to high confidence in the prediction.
  • Complementary: $a(q,c)\neq a(q,\phi)$ but $a=a(q,c)\cup a(q,\phi)$; neither source is sufficient alone, but their combined information yields the final answer.
  • Conflicting: $a(q,c)\neq a(q,\phi)$ and $(a=a(q,c))\oplus(a=a(q,\phi))$; the final answer aligns exclusively with one knowledge source while rejecting the other.
  • Irrelevant: $c\perp q$ such that $a(q,c)$ is null or nonsensical; the context $c$ contains no semantic information relevant to the query $q$. In this case, the model ignores the context entirely and relies solely on PK ($a=a(q,\phi)$) to formulate a response.
  • Knowledge Suppression (noise): $a(q,c)=a(q,\phi)\neq a$; both sources agree on an answer, yet the model fails to utilize either, producing an erroneous or unrelated prediction.
This taxonomy enables a systematic investigation of how PK and CK interact throughout NLE generation, moving beyond a binary conflict to capture a spectrum of dynamics. We reformulate this intent-driven answer control and $w_{b}$ denote the joint intent of $w_{c}$ and $w_{p}$ that illustrates the intrinsic model behaviour in generating the final answer $a$, considering both the PK and CK governed by the underlying interaction scenario. Unlike in Minder et al. (2025), where $w_{c}$ and $w_{p}$ are represented using one orthonormal direction in a learned rank-1 projection subspace, we represent $w_{c}$ and $w_{p}$ using two orthonormal directions, and $w_{b}\rightarrow w_{c}$ and $w_{b}\rightarrow w_{p}$ indicate the individual contributions of CK and PK, respectively, in generating $a$. Then the joint intent $w_{b}=f(w_{c},w_{p})$ indicates different knowledge interaction scenarios decided by individual contributions from $w_{c}$ and $w_{p}$. We learn this knowledge interaction function using a rank-2 projection subspace. The prompt template for the three intent-driven answer control is described in Table 1 in §A.1.2.
@@ 3.3 Localize Intent-Guided PK-CK Interactions
Identifying important layers for rank-1 projection subspace. We follow the activation patching-based mechanistic interpretability approach, Patchscope (Minder et al., 2025). They construct two minimally different prompts $s$ and $t$ as ‘source’ and ‘target’, with the same $q$ and $c$, only differing by the intent, resulting in two different intended answers (following the prompt template in Table 1). To identify the layers capturing the intent present in $s$, during the forward pass of the target $p(.|t)$, the activation from the hidden state of a particular layer is replaced by the activations from the same layer with the source $p(.|s)$, resulting in $\tilde{p}(.|t)\approx p(.|s)$. Since they only consider the ‘conflicting’ behaviours in understanding the knowledge interaction in LMs, they construct two patching datasets capturing opposite directions: $\mathcal{D}_{w}^{(c\rightarrow p)}=\{((q,c,w_{c}),a(q,c)),\;((q,c,w_{p}),a(q,\phi))\}$ for encoding the CK direction, and $\mathcal{D}_{w}^{(p\rightarrow c)}=\{((q,c,w_{p}),a(q,\phi)),\;((q,c,w_{c}),a(q,c))\}$ for encoding the PK direction. Layers capturing CK and PK directions, respectively, are selected as:
@@
$\mathbb{L}_{p\rightarrow c}$ $=\{\,l\in[1,L]\mid\tilde{p}(a(q,\phi)\mid q,c,w_{c})\geq\tau_{p}\,\},$
where $\tau_{c}$, $\tau_{p}$ are hyperparameters. For more details, please refer to Minder et al. (2025).
Identifying important layers for rank-2 projection subspace (showing the individual knowledge contribution towards generating $a$, i.e, $w_{b}\rightarrow w_{c}$ and $w_{b}\rightarrow w_{p}$). We again follow Patchscope and prepare two patching datasets $\mathcal{D}_{w}^{(b\rightarrow p)}=\{((q,c,w_{b}),a),\;((q,c,w_{p}),a(q,\phi))\}$ for encoding the contribution of PK interacting with CK in generating $a$, and $\mathcal{D}_{w}^{(b\rightarrow c)}=\{((q,c,w_{b}),a),\;((q,c,w_{c}),a(q,c))\}$ for encoding the contribution of CK interacting with PK in generating $a$. Important layers $\mathbb{L}_{b\rightarrow c}$ and $\mathbb{L}_{b\rightarrow p}$ are selected using Equation 2 ( (hyperparameter details are described in Figure 10 in §A.1.3)).
@@
Theorem 3.1 (Non-identifiability under rank-1).
@@
Let $a=\langle\vec{\mathbf{v}},\vec{\mathbf{u}}_{CK}\rangle$ and $b=\langle\vec{\mathbf{v}},\vec{\mathbf{u}}_{PK}\rangle$. For any $(c,p)$, choose $(c^{\prime},p^{\prime})=(c+\delta,\,p-\tfrac{a}{b}\delta)$ with $\delta\neq 0$. Then $ac+bp=ac^{\prime}+bp^{\prime}$, so infinitely many $(c^{\prime},p^{\prime})$ yield the same observation. Thus the mapping $(c,p)\mapsto\alpha$ is non-injective, and the individual contributions cannot be disentangled. ∎
Encoding the intent via rank-2 subspace projection. Once we identify important layers for $w_{b}\rightarrow w_{c}$ and $w_{b}\rightarrow w_{p}$, we learn the joint intent function $f$ encoding how individual knowledge contributions $w_{b}\rightarrow w_{c}$ and $w_{b}\rightarrow w_{p}$ are mixed to generate the final answer $a$ using the common layers from $\mathbb{L}_{b\rightarrow c}$ and $\mathbb{L}_{b\rightarrow p}$ via a rank-2 projection subspace $\mathbf{P}$ spanned by two orthogonal directions $\vec{\mathbf{u}}\in\mathbb{R}^{d*2}$ as $\mathbf{P}=\vec{\mathbf{u}}(\vec{\mathbf{u}}^{T}\vec{\mathbf{u}})^{-1}\vec{\mathbf{u}}^{T}$. Since we aim to identify individual contributions from the individual directions, following Minder et al. (2025), we consider those two basis vectors as orthonormal and hence the rank-2 projection subspace $\mathbf{P}$ is reduced to $\mathbf{P}=\vec{\mathbf{u}}\vec{\mathbf{u}}^{T}$, since $\vec{\mathbf{u}}^{T}\vec{\mathbf{u}}=\mathbf{I_{2}}$. Considering $\vec{\mathbf{u}}=[\vec{\mathbf{u_{c}}};\vec{\mathbf{u_{p}}}]$, at any sequence step $i$, the hidden representation $\vec{\mathbf{h_{i}}}\in\mathbb{R}^{d}$ for the token $e_{i}$ in the NLE $\mathcal{E}$ can be linearly decomposed as:
@@
Learning the rank-2 projection subspace. Let $\tilde{\vec{\mathbf{h}}}^{w_{p}}=(\mathbf{I}-\mathbf{P})\,\vec{\mathbf{h}}+\vec{\mathbf{u}}_{c}\,\langle\vec{\mathbf{u}}_{c}^{T},\vec{\mathbf{h}}\rangle+\vec{\mathbf{u}}_{p}\,\langle\vec{\mathbf{u}}_{p}^{T},\vec{\mathbf{h}}^{\,p}\rangle$, and $\tilde{\vec{\mathbf{h}}}^{w_{c}}=(\mathbf{I}-\mathbf{P})\,\vec{\mathbf{h}}+\vec{\mathbf{u}}_{c}\,\langle\vec{\mathbf{u}}_{c}^{T},\vec{\mathbf{h}}^{\,c}\rangle+\vec{\mathbf{u}}_{p}\,\langle\vec{\mathbf{u}}_{p}^{T},\vec{\mathbf{h}}\rangle$ be two patched decompositions of the last-token hidden representation of the NLE $\mathcal{E}$ by replacing the activation from a particular layer $l\in\mathbb{L}_{b\rightarrow p}\cup\mathbb{L}_{b\rightarrow c}$ during the forward pass of $p(\cdot\mid q,c,w_{b})$ with the activations from the same layer of the forward pass of $p(\cdot\mid q,c,w_{p})$ and $p(\cdot\mid q,c,w_{c})$, respectively. Then we learn the projection subspace $\mathbf{P}$ spanned by the two orthonormal basis vectors $\vec{\mathbf{u}}=[\vec{\mathbf{u}}_{c};\vec{\mathbf{u}}_{p}]$ by minimizing the two objectives sequentially, maintaining orthonormal directions between $\vec{\mathbf{u}}_{c}$ and $\vec{\mathbf{u}}_{p}$:
$\mathcal{J}_{p}$ $=-\frac{1}{N}\sum_{n=1}^{N}\log\bigl(\tilde{p}\!\left(a(q_{n},\phi)\mid\tilde{\vec{\mathbf{h}}}_{(n)}^{w_{p}}\right)\bigr),$ (8)
$\mathcal{J}_{c}$ $=-\frac{1}{N}\sum_{n=1}^{N}\log\bigl(\tilde{p}\!\left(a(q_{n},c_{n})\mid\tilde{\vec{\mathbf{h}}}_{(n)}^{w_{c}}\right)\bigr).$ (9)
Normaized PK-CK score. At any sequence step $i$, given the the hidden representation $\vec{\mathbf{h_{i}}}\in\mathbb{R}^{d}$ for the token $e_{i}$ in the NLE $\mathcal{E}$, let $c_{i}=\lvert\langle\vec{\mathbf{u_{c}}}^{T},\vec{\mathbf{h_{i}}}\rangle\lvert$ and $p_{i}=\lvert\langle\vec{\mathbf{u_{p}}}^{T},\vec{\mathbf{h_{i}}}\rangle\lvert$, then the normalized contribution from CK and PK can be computed as ${\alpha_{i}}^{c}=c_{i}/(c_{i}+p_{i})$ and ${\alpha_{i}}^{p}=p_{i}/(c_{i}+p_{i})$, and ${\alpha_{i}}^{c},{\alpha_{i}}^{p}\in[0,1]$ satisfying the identifiability of individual knowledge contribution (Theorem 3.1) under rank-2 subspace. The methodology of assigning the PK-CK direction to the basis vector $\vec{\mathbf{u}}$ is described in §A.1.1.